# EVINGCA: Adaptive Graph Clustering with Evolving Neighborhood Statistics

### v1

@@ VI-I Runtime Analysis
For small datasets ($N<5{\times}10^{3}$), EVINGCA completes within 0.10.10.20.2 s, showing approximately linear growth with $N$. Between the Spiral (312 points, 0.02 s) and Labirynth (3546 points, 0.20 s) datasets, the runtime increases roughly $10\times$ for an $11\times$ increase in $N$, consistent with the expected $O(N\log N)$ scaling. At $N=10^{4}$, runtimes remain below 0.5 s for moderate dimensions (e.g., Pendigits, 16D), confirming the near-linear trend predicted by Section IV-B.
@@
EVINGCA consistently ranks between centroid-based methods (K-means, GMM) and graph- or density-based algorithms (HDBSCAN, Spectral) in raw speed. Across the ten datasets shown in Table II:
@@
  • On medium-size datasets (5k–10k samples, $D\leq 100$), EVINGCA averages $\approx 0.4$ s versus $\approx 1$ s for HDBSCAN and >3>3 s for Spectral.
  • In high-dimensional cases ($D\geq 500$), EVINGCA outperforms HDBSCAN by one to two orders of magnitude (e.g., 0.76 s vs. 25.6 s on HAR, 2.1 s vs. 62.6 s on Fashion-MNIST).
K-means remains the absolute fastest method due to its minimal per-iteration cost and absence of neighbor queries, but EVINGCA achieves significantly greater expressivity with only a modest runtime penalty.

### v2

@@ VI-I Runtime Analysis
For datasets with 2–16 dimensions, EVINGCA completes within 0.010.010.40.4 s and shows approximately linear growth with $N$. For example, between the Spiral (312 points, 0.02 s) and Labirynth (3546 points, 0.20 s) datasets, the runtime increases roughly $10\times$ for an $11\times$ increase in $N$. Then on Pendigits (10000 points, $0.4$ s), runtime increases by $2\times$ for a $3\times$ increase in $N$.
@@
EVINGCA is slower than K-means and faster than Spectral. On low-dimensional datasets, HDBSCAN often runs faster, however as dimensionality increases, EVINGCA overtakes it and in some cases, outpaces GMM as well.
@@
[…]
  • K-means remains the fastest method in the majority of cases, although it is occasionally superseded by HDBSCAN on low-dimensional datasets.
Since part of EVINGCA’s overhead arises from its current Python-based implementation, this comparison is likely conservative. Further runtime reductions are expected from a fully vectorized or Cython-optimized variant.